32. A stock has a beta of 1.8, the expected return on the market is 10...

32.

A stock has a beta of 1.8, the expected return on the market is 10 percent, and the risk-free rate is 6.5 percent. What must the expected return on this stock be?


rev: 09_20_2012

13.31%

12.8%

12.16%

24.5%

13.44%

34.

The Down and Out Co. just issued a dividend of $2.51 per share on its common stock. The company is expected to maintain a constant 4 percent growth rate in its dividends indefinitely. If the stock sells for $40 a share, what is the company's cost of equity? (Do not round your intermediate calculations.)


rev: 09_20_2012

10.53%

10%

10.28%

6.63%

11.05%

35.

You own a stock portfolio invested 25 percent in Stock Q, 20 percent in Stock R, 5 percent in Stock S, and 50 percent in Stock T. The betas for these four stocks are 1.37, 1.76, 1.75, and 1.51, respectively. What is the portfolio beta?


rev: 09_20_2012

1.57

1.51

1.61

1.46

1.54
Answer #1

1.Expected return=risk-free rate +Beta*(market rate- risk-free rate)

=6.5+1.8*(10-6.5)

=12.8%

2. cost of equity=(D1/Current price)+Growth rate

=(2.51*1.04)/40+0.04

=10.53%(Approx).

3.Portfolio beta=Respective beta*Respective investment weight

=(0.25*1.37)+(0.2*1.76)+(0.05*1.75)+(0.5*1.51)

=1.54(Approx).